Solve the system of equations by the addition method.
x/4-y/12=-1
x/3+y/9=-2

Answer by mananth(16949)

x/4-y/12=-1
LCD =12
3x-y /12 =-1multiply by 12
3x-y=-12.........................1
x/3+y/9=-2
lCD=9
(3x+y)/9 = -2
multiply by 9
3x+y=-18.........................2
Add (1)&(2)
6x=-12+-18
6x=-30
/6
x=-5
...
plug value of x in 2
-15+y=-18
+15 +15
y=-3
...
m.ananth@hotmail.ca